A line through the points (9,2) and (7,x) has a slope of 7/3. Find x.
I think this is how to start. (7-2/7-x)=7/3
(7/7-x)=7/3
7-x-=7
-7 -7
_________
-x=0
-- --
- 1 -1
x=o

Found 2 solutions by lwsshak3, Alan3354:
Answer by lwsshak3(11628)

A line through the points (9,2) and (7,x) has a slope of 7/3. Find x
**
slope=∆y/∆x=change in y/change in x
(2-x)/(9-7)=7/3
(2-x)/2=7/3
6-3x=14
3x=-8
x=-8/3

Answer by Alan3354(69443)

A line through the points (9,2) and (7,x) has a slope of 7/3. Find x.
I think this is how to start. (7-2/7-x)=7/3
(7/7-x)=7/3
7-x-=7
-7 -7
_________
-x=0
-- --
- 1 -1
x=o
==================
A line through the points (9,2) and (7,x) has a slope of 7/3. Find x.
---------
Not a well posed problem, it has x in the y position.
If it said:
"A line through the points (9,2) and (7,y) has a slope of 7/3. Find y." it would make more sense.
-------------
Slope m = diffy/diffx = (2-y)/(9-7) = 7/3
(2-y)/(9-7) = 7/3
Cross multiply
3(2-y) = 2*7
6-3y = 14
-3y = 8
y = -8/3